OpenStudy (natalie1234):

PLEASE HELP WILL GIVE MEDAL!!(: Find the greatest common factor: 20x^6y + 40x^4y^2 − 10x^5y^5

OpenStudy (jdoe0001):

\(\large 20x^6y + 40x^4y^2 - 10x^5y^5\implies \begin{array}{llll} 20x^6y \implies & 10\cdot 2\cdot x^4\cdot x^2\cdot y \\ \quad \\ 40x^4y^2\implies &10\cdot 4\cdot x^4\cdot y\cdot y \\ \quad \\ 10x^5y^5\implies & 10\cdot x^4\cdot x\cdot y^4\cdot y \end{array}\) what do you think?